a man is six times as old as his son. twenty years later, the man will be twice as old as his son. find the age of the man when his was born
Answer by ankor@dixie-net.com(22740)

Write an equation for each statement
:
a man is six times as old as his son.
m = 6s
:
twenty years later, the man will be twice as old as his son.
m + 20 = 2(x+20)
m + 20 = 2s + 40
m = 2s + 40 - 20
m = 2s + 20
replace m with 6s (from the 1st statement)
6s = 2s + 20
6s - 2s = 20
4s = 20
s = 20/4
s = 5 yrs old is the son
then
6(5) = 30 yrs is the man's age
:
"Find the age of the man when his son was born"
30 - 5 = 25 yrs